please answer me one more question if its possible. my new problem is
FIND THE HEIGHT AND BASE OF A TRIANGLE IF THEIR SUM IS 20 AND THEIR AREA IS A MAXIMUM.

i rlly dont get maximum and minimum values. what does that mean? and whats the difference when it says minimum or maximum. how can i set it up?
thanx truly

Answer by checkley71(8403)

Area of a triangle=h*b/2
b+h=20
if b=h then:
10*10/2=10*5=50 is the maximum area.
proof:
use 9+11 20
9*11/2=99/2=45.5
11*9/2=99/2=45.5